Calculate how much chromium (as grams) can be electrodeposited in a one day when the current is 1.50 A. Solution contains Cr3+ ions.
1
Expert's answer
2012-04-19T07:39:05-0400
m = Mw*I*t/n*F

m - mass
Mw - molecular weight
n - the number of electrons involved in the process
I - electric current
F - Faraday constant
t - time

m = 51.996*1.5* (24 * 60 *60)/3*96 485.3 = 23.3 g